# LogicGame: Benchmarking Rule-Based Reasoning Abilities of Large Language Models

### v1

@@ Abstract
22footnotetext: Work done when JG and JC interned at Zhipu AI.
@@ 4.2 Main Results and Analysis
The average performance of 11 LLMs on LogicGame across our Chinese and English bilingual dataset, measured by AP-Acc, is presented in Table 2. The results indicate that reasoning remains a significant challenge for LLMs. Even the strongest model can’t reach an overall score of 20%. As the difficulty level increases, the performance of LLMs drops significantly, with many models dropping to 0% score at Level 3, especially in Planning tasks. This suggests that even advanced LLMs struggle with complex reasoning tasks.
@@
Interestingly, the performance of models in Execution and Planning tasks varies. Some top-performing models, such as claude-3.5-sonnet, demonstrates superior performance in Execution compared to Planning (22.51% vs. 15.80%). Conversely, llama-3-70b-chat excels in Planning over Execution (11.00% vs. 6.53%). Even within model families, the relative performance differences are evident. In the GPT family, gpt-4-turbo-0409 outperforms gpt-4o in Planning tasks (15.32% vs. 13.21%), while gpt-4o shows better performance in Execution tasks (19.73% vs. 15.89%).
@@
claude-3.5-sonnet 55.56 17.78 11.11 5.56 22.51 42.12 15.77 1.79 3.51 15.80 19.15
gpt-4o 43.89 16.87 12.48 5.69 19.73 32.74 16.41 3.70 0.00 13.21 16.48
gpt-4-turbo-0409 38.89 14.45 6.82 3.41 15.89 43.85 13.98 3.45 0.00 15.32 15.61
qwen2-72b-instruct 31.11 10.00 3.39 0.00 11.13 32.81 10.91 0.00 0.00 10.93 11.03
llama-3-70b-chat 18.61 7.50 0.00 0.00 6.53 32.91 11.07 0.00 0.00 11.00 8.77
claude-3-haiku 14.52 6.67 4.45 0.00 6.41 19.28 3.51 0.00 0.00 5.70 6.06
glm-4-9b 18.57 1.28 0.00 0.00 4.96 18.26 3.57 0.00 0.00 5.46 5.22
internlm-2.5-7b-chat 6.98 2.47 1.28 0.00 2.69 18.52 0.00 1.85 0.00 5.10 3.90
qwen2-7b-instruct 10.58 2.22 0.00 0.00 3.20 14.62 1.85 0.00 0.00 4.12 3.66
llama-3-8b-chat 2.94 0.00 0.00 0.00 0.74 9.38 2.78 0.00 0.00 3.04 1.89
mistral-7b-instruct 4.16 0.00 0.00 0.00 1.04 5.77 0.00 0.00 0.00 1.45 1.25
Table 2: Performance of 11 models on LogicGame measured by AP-Acc. The highest performance is bold.

### v2

@@ Abstract
22footnotetext: Work done when JG and JC interned at Zhipu AI.
@@ 4.2 Main Results and Analysis
The average performance of 11 LLMs on LogicGame across our Chinese and English bilingual dataset, measured by AP-Acc, is presented in Table 2. The results indicate that reasoning remains a significant challenge for LLMs. Even the strongest model can’t reach an overall score of 21%. As the difficulty level increases, the performance of LLMs drops significantly, with many models dropping to 0% score at Level 3, especially in Planning tasks. This suggests that even advanced LLMs struggle with complex reasoning tasks.
@@
Interestingly, the performance of models in Execution and Planning tasks varies. Some top-performing models, such as claude-3.5-sonnet, demonstrates superior performance in Execution compared to Planning (22.50% vs. 16.67%). Conversely, llama-3-70b-chat excels in Planning over Execution (10.19% vs. 5.56%). Even within model families, the relative performance differences are evident. In the GPT family, gpt-4-turbo-0409 outperforms gpt-4o in Planning tasks (16.21% vs. 13.43%), while gpt-4o shows better performance in Execution tasks (19.44% vs. 15.83%).
@@
claude-3.5-sonnet 55.56 17.78 11.11 5.56 22.50 44.44 16.67 1.85 3.70 16.67 20.93
gpt-4o 43.33 16.67 12.22 5.56 19.44 33.33 16.67 3.70 0.00 13.43 18.43
gpt-4-turbo-0409 38.89 14.44 6.67 3.33 15.83 46.30 14.81 3.70 0.00 16.21 17.04
glm-4-plus 34.44 11.11 7.78 1.11 13.61 35.19 16.67 1.85 0.00 13.43 14.64
qwen2-72b-instruct 31.11 10.00 3.33 0.00 11.11 33.33 11.11 0.00 0.00 10.93 12.01
llama-3-70b-chat 16.67 5.56 0.00 0.00 5.56 31.48 9.26 0.00 0.00 10.19 8.28
claude-3-haiku 14.44 6.67 4.44 0.00 6.39 20.37 3.70 0.00 0.00 6.02 6.80
glm-4-9b 18.89 1.11 0.00 0.00 4.72 18.52 3.70 0.00 0.00 5.56 5.76
internlm-2.5-7b-chat 6.67 2.22 1.11 0.00 2.50 18.52 0.00 1.85 0.00 5.10 4.28
qwen2-7b-instruct 10.00 2.22 0.00 0.00 3.06 14.81 1.85 0.00 0.00 4.17 3.29
mistral-7b-instruct 3.33 0.00 0.00 0.00 0.83 5.56 0.00 0.00 0.00 1.39 1.97
llama-3-8b-chat 1.11 0.00 0.00 0.00 0.28 5.56 1.85 0.00 0.00 1.85 0.81
Table 2: Performance of 11 models on LogicGame measured by AP-Acc as the arithmetic mean of both en and zh version. The highest performance is bold.